OpenStudy (anonymous):

Use the matrix method to solve: x + 4y = 8 2x + y = 9 x=? y=?

OpenStudy (anonymous):

x=3.43 y=1.14

OpenStudy (anonymous):

\[\left(\begin{array}{cc|c}1 & 4 & 8 \\ 2 & 1 & 9\end{array}\right)\] Substract 2*row 1 from row 2: \[\left(\begin{array}{cc|c}1 & 4 & 8 \\ 0 & -7 & -7\end{array}\right)\] now multiply row 2 by -1/7: \[\left(\begin{array}{cc|c}1 & 4 & 8 \\ 0 & 1 & 1\end{array}\right)\] substract 4 * row 2 from row 1: \[\left(\begin{array}{cc|c}1 & 0 & 4 \\ 0 & 1 & 1\end{array}\right)\] x = 4, y = 1